# Challenges in applying the magnetic field circulation in Ampère-Maxwell’s law

### v1

@@ Abstract
We investigate how students in introductory electromagnetism courses apply the idea behind the magnetic field circulation in the context of the Ampère-Maxwell law. For this purpose, we designed a metacognitive pencil-and-paper question, presented it to 65 students, and analysed their responses using phenomenography. We complemented our research by conducting interviews with 12 students. The results show that students tend to think that they can only use the magnetic field circulation to calculate magnetic fields if the curve used is symmetric, and that they do not take into account the shape of the magnetic field lines when applying it. We also find that some students believe that a variable electric field must cross the entire Amperian curve in order to apply Ampère-Maxwell’s law and find the magnetic field.
@@ 1 Introduction
Maxwell’s equations are the cornerstone of classical electromagnetic theory. In the realm of Physics Education Research (PER), there have been many studies of the conceptual difficulties that students face in applying them, with particular attention paid to Gauss’s law of the electric field [1, 2, 3, 4, 5, 6, 7, 8] and Faraday’s law [9, 10, 11, 12, 13], but less research on Ampère’s law and, in particular, Ampère-Maxwell’s law.
With regard to Ampère’s law, a significant percentage of students use reasoning based on memorisation and mechanical application of the information received in class and incorrectly consider that the magnetic field circulation can always be expressed as the product of the magnetic field and the length of the curve. They also think that the magnetic field used in the calculation of the line integral originates only from the currents crossing a surface bounded by the curve [1, 2]. Additionally, they assume that when the net current null, the magnetic field also vanishes [1, 2, 14]. Students often do not use the available information about the magnetic field to calculate the circulation, nor do they realise that the magnetic field circulation is a sum of elements $\vec{B}\cdot\overrightarrow{dl}$ along a closed curve [14]. They also confuse the concepts of circulation with magnetic field and circulation with flux [7, 8].
@@
This work is a continuation of research into students’ conceptual difficulties with Ampère-Maxwell’s law and is part of a wider project with the ultimate aim of developing a teaching-learning sequence. In this context, and considering the conceptual difficulties identified in the previous study, we have extended our original research to delve deeper into how students apply the magnetic field circulation in situations where they have to calculate it explicitly from the magnetic field. Our research question is:
@@
  • How do students apply the magnetic field circulation in the context of Ampère-Maxwell’s law?
@@
To address this research question, we designed a metacognitive pencil-and-paper question that led students to think about their own learning. This type of question has been widely recommended to investigate learning by understanding and active learning [16, 17]. The question presents a problem and a possible solution, so that the students must evaluate whether the solution is correct or not. To deepen the explanations received, we conducted semi-structured interviews with several groups of students. In the following sections, we describe the methodology used. We then present the results obtained, followed by a discussion and, finally, conclusions and implications for the teaching of Ampère-Maxwell’s law.
2 Experimental design
@@
The question posed to the students shows a uniform electric field with magnitude increasing with time, confined to a cylindrical region of radius $R$, and states that the magnetic field is to be determined at a point $P$ located at a certain distance from the symmetry axis of the cylindrical region. The step-by-step solution proposed by a student to calculate the magnetic field as a function of the parameters of the problem is shown below. The proposed solution is correct, but it has the peculiarity that the closed curve chosen by the student exhibits a semicircle shape.
@@
    To calculate the magnetic field circulation, it is necessary to know the shape of the field lines in the context of the defined curve $C$, since its definition is a sum of elements $\vec{B}\cdot\overrightarrow{dl}$.The Ampère-Maxwell law correlates the magnetic field circulation along the curve $C$ with the net current crossing a surface bounding it, without implying cause-effect relationships [18, 19].
3 Methodology
The question was answered by 65 undergraduate physics students in introductory electromagnetism courses at the University of the Basque Country (UPV/EHU) and the University of the Republic of Uruguay (UDELAR). These courses were taught by experienced university professors using standard teaching methods. The course duration was similar in both universities, lasting 15 weeks, with 4 hours of theory classes and an additional 2 hours per week dedicated to problem solving. The course content was based on the most recommended textbooks for introductory electromagnetism courses in the United States, Latin America and Spain [20].
The question was proposed to the students after they had received instruction on Ampère-Maxwell’s law and was analysed by two of the researchers using phenomenographic methodology [21]. This method made it possible to identify different categories of explanations that reflected the differences in students’ understanding. At the beginning of the analysis, each researcher evaluated 20 written tests. A consensus was then reached on the categories of understanding and explanation that emerged from the students’ responses. Each researcher then analysed the remaining tests individually. At the end of this process, the results obtained were compared, and the reliability of the analysis was assessed using Cohen’s kappa coefficient, which yielded a value of 0.95, indicating significant agreement [22].
To complement the study, we conducted semi-structured group interviews using the same question with three groups of two students and two groups of three students. Each interviews lasted approximately 20 minutes and in all the cases were conducted by the first author of this paper. The interviewer asked the students to discuss the question aloud and encouraged them to provide detailed explanations of their understanding using non-directive questions such as What do you mean by that?, Could you explain more?, Is there anything else you’d like to add about this step? All interviews were audio-recorded and transcribed. The extracts presented in the results section illustrate the sequence of interpretations and reasoning that characterise the processes of students’ development of situated cognition through problem solving. Fictitious names have been used in the presentation of results to ensure anonymity.
@@ 4 Results
From the phenomenographic analysis of students’ written responses, several explanatory categories emerge about how students reason when they engage in metacognitive reflection about solving the problem (see table 1).
Table 1: Answer categories and their percentages for question.
Category Description of the category %
A They agree with the analysis. 15.4
B They do not agree with the calculation of the magnetic field circulation. 15.4
C They do not agree with the choice of the curve. 20.0
D Other/Incoherent. 15.4
E No answer. 33.8
In category A we include students who justify their agreement with the solution presented. Let’s look at a representative example of this category:
@@
“Yes, because in $C_{1}$, $\vec{B}\cdot\overrightarrow{dl}=0$ since they are orthogonal throughout $C_{1}$, and in $C_{2}$, $\vec{B}\cdot$ $\overrightarrow{dl}=Bdl$ since they are collinear.” (UDELAR-7, see Fig. 1)
Figure 2: Hand-drawn sketch given by student UDELAR-7.
@@
Student UDELAR-7 divides the closed curve into two sections, $C_{1}$ and $C_{2}$, and correctly recognises that the contribution to the magnetic field circulation along $C_{1}$ is null. This indicates that, although they do not directly represent the magnetic field along $C_{1}$, they understand that the magnetic field is perpendicular to each element of the curve. On the other hand, the student observes that the magnetic field is collinear with section $C_{2}$. In this way, they demonstrate an adequate understanding of the shape of the magnetic field lines and correctly interpret and apply the concept of magnetic field circulation.
In category B we include answers that disagree with the calculation of the magnetic field circulation because they fail to recognise that the contribution of the magnetic field along the straight section of the curve is null. Let’s look at examples of category B:
“dl is $(2+\pi)a$ and the result is $B=\frac{\mu_{0}\varepsilon_{0}}{(2+\pi)a}\frac{d\Phi_{\mathrm{E}}}{dt}$.”(UPV-28)
@@
“The induced field B is perpendicular to the field E at every point and its lines are closed. It is not perpendicular to the straight section of curve C, so it should have considered that section in the integral: $B(\pi a+2a)$.” (UDELAR-4)
@@
Student UPV-28 believes that the magnetic field circulation should be calculated as the product of the magnetic field and the length of curve $C$. However, they do not recognise that the magnetic field along the straight section of curve $C$ does not contribute to the magnetic field circulation. On the other hand, student UDELAR-4 considers the direction of the magnetic field, acknowledging its perpendicularity to the electric field, but struggles to use this information effectively to calculate the magnetic field circulation, probably due to an incomplete understanding of circulation.
@@
In category C we group the answers that express disagreement with the choice of curve $C$ for the calculation of the magnetic field and support their arguments based on symmetry considerations. Let’s look at a few examples of category C:
“No, because in my case the integration curve would be $2\pi a$ and not $\pi a$.” (UPV-30)
“I think it is wrong, because the whole surface is not affected by the electric field, you have to consider the whole circle.” (UPV-5)
From student UPV-30’s answer, we can deduce that their disagreement with the solution of the question is due to the symmetry of the curve used, considering that a complete circle should be used. On the other hand, the UPV-5 student argues that a complete circle should be used so that the entire variable electric field crosses the surface bounded by the curve.
@@
As we can see, Lucía clearly explains her reasoning. She states that the calculation of the magnetic field is only possible if the closed curve used is symmetrical.
In another interview, we identified an exchange between two students that is in line with what was expressed by the UPV-5 student. Below we present an extract from the interview in which the students exchange ideas about the shape of the closed curve.
@@ 5 Discussion of results
When analysing the students’ answers to the research question, we found that around 15% adequately justify why they agree with all the steps in solving the question, showing an understanding of the application of magnetic field circulation in the context of Ampère-Maxwell’s law. We also observed that a similar percentage disagrees with the calculation of the magnetic field circulation, because they do not recognise that the magnetic field does not contribute to the circulation in the straight section. In this sense, we identified reasoning that could be related to the rote application of Ampère-Maxwell’s law, where students believe that the magnetic field circulation can always be expressed as the product between the magnetic field and the length of the closed curve. This type of reasoning has been documented in the literature in relation to Ampère’s law in different contexts [1, 2, 14].
@@
Furthermore, we found that in some cases students do not consider the shape of the magnetic field lines when applying Ampère-Maxwell’s law, while in others they do not use the magnetic field information adequately to calculate the magnetic field circulation, nor do they recognise that the magnetic field circulation is a sum (see for example the first interview in section 4). These findings converge with those reported by Wallace and Chasteen [14] in the context of Ampère’s law.
@@
We found that one in five students disagree with the choice of curve for calculating the circulation. Many of them believe that Ampère-Maxwell’s law can only be used to calculate magnetic fields if the curve used is symmetrical, which could be due to a functional fixation [23] associated with the most used procedure in class, which argues for symmetry to calculate magnetic fields using Ampère-Maxwell’s law. In addition, we found students who disagree with the shape of the curve used in the problem, considering that the variable electric field must cross the entire Amperian curve in order to be able to apply Ampère-Maxwell’s law and calculate the magnetic field. Our results suggest that students are unaware that Ampère-Maxwell’s law correlates the magnetic field circulation along a closed curve with the net current that crosses a surface bounding it.
@@
This conceptual difficulty could be due to the belief that Ampère-Maxwell’s law establishes cause-effect relationships between the terms involved, in particular with the idea that currents (conduction and displacement) crossing a surface bounded by a curve must be the cause of the magnetic field computed in the line integral. This could lead them to argue that the curve must enclose the entire electric field. Another possible explanation is related to the belief that the magnetic field used in Ampère’s law is caused by enclosed currents [1, 2] and its origin could be in the way Maxwell’s equations are presented in most introductory physics textbooks, where they are described as cause-effect relationships between their different terms [20].

### v2

# Understanding student challenges with circulation and Ampère-Maxwell law

@@ Abstract
We investigated how students apply the concept of circulation when faced with problems related to the Ampere-Maxwell law. For this purpose, we designed a metacognitive pencil-and-paper question, presented it to 65 students, and analysed their responses using phenomenography. We complemented our research by conducting interviews with 12 students. The results show that students tend to think that they can only use the magnetic field circulation to calculate magnetic fields if the curve used is symmetric, and that they do not take into account the shape of the magnetic field lines when applying it. We also find that some students believe that a variable electric field must cross the entire Amperian curve in order to apply Ampère-Maxwell’s law and find the magnetic field.
@@ 1 Introduction
Maxwell’s equations are the cornerstone of classical electromagnetic theory. In Physics Education Research (PER), there have been many studies of the conceptual difficulties that students face in applying them, with particular attention paid to Gauss’s law of the electric field [1, 2, 3, 4, 5, 6, 7, 8] and Faraday’s law [9, 10, 11, 12, 13], but less research on Ampère’s law and, in particular, Ampère-Maxwell’s law.
With regard to Ampère’s law, a significant percentage of students use reasoning based on memorisation and mechanical application of the information received in class and incorrectly consider that the magnetic field circulation can always be expressed as the product of the magnetic field and the length of the curve. They also think that the magnetic field used in the calculation of the line integral originates only from the currents crossing a surface bounded by the curve [1, 2]. Additionally, they assume that when the net current is null, the magnetic field also vanishes [1, 2, 14]. Students often do not use the available information about the magnetic field to calculate the circulation, nor do they realise that the magnetic field circulation is a sum of elements $\vec{B}\cdot\overrightarrow{dl}$ along a closed curve [14]. They also confuse the concepts of circulation with magnetic field and circulation with flux [7, 8].
@@
This work is a continuation of research into students’ conceptual difficulties with Ampère-Maxwell’s law [15] and is part of a wider project with the ultimate aim of developing a teaching-learning sequence [15, 16]. In this context, and considering the conceptual difficulties identified in the previous study [15], we have extended our original research to further explore the various qualitative approaches students use when applying magnetic field circulation in situations requiring explicit calculation from the magnetic field. Our research question is:
@@
  • In what distinct qualitative ways do students apply magnetic field circulation in the context of Ampère-Maxwell’s law?
@@
To address this research question, we designed a metacognitive pencil-and-paper question that led students to think about their own learning. This question proposes a problem and a possible solution, so that the students must evaluate whether the solution is correct or not. To deepen the explanations received, we conducted semi-structured interviews with several groups of students.
Metacognition refers to an individual’s ability to be aware of and regulate their own thought processes. It encourages individuals to reflect on their thinking, critically analyzing it for deeper understanding and continuous improvement [17]. While metacognition is often defined as “thinking about thinking” or “learning about learning,” it also involves the use of problem-solving strategies such as planning, monitoring, and evaluating [17]. This approach has been widely recommended for investigating learning through understanding and active engagement, as it is broadly recognized that developing metacognitive skills is a crucial component of the learning process [18, 19]. In the following sections, we outline the methodological approach and the context of the research and design. We then present the results obtained, followed by a discussion, and conclude with the implications for teaching Ampère-Maxwell’s law.
2 Methodological approach
Students’ conceptions have been investigated through different tasks and in different contexts, and the coherence of these conceptions is a key aspect to consider [20]. This issue raises the need for studies that describe the variability in students’ conceptions. Phenomenography has been proposed and employed as a methodological approach to analyse and explain this variability [21, 22]. Here, we adopt a phenomenographic approach to explore the various qualitative ways in which individuals experience, conceptualize, perceive, and understand different phenomena and aspects of the world around them [21]. Marton and Booth explain that
“in phenomenography individuals are seen as the bearers of different ways of experiencing a phenomenon and as bearers of fragments of differing ways of experiencing that phenomenon.” [22, p 114].
In this approach, the description elicited from learners is collective, leaving aside individual voices. Phenomenography examines how diverse ways of perceiving and understanding reality (concepts and associated modes of reasoning) can be organised into categories that describe reality. These categories represent shared patterns of understanding, visible across many individuals, and thus reflect a kind of collective knowledge:
“The same description categories appear in different situations. The set of categories is thus stable and can be applied, even if individuals “move” from one category to another on different occasions.” [21, p 195].
Marton and Booth [22] propose certain criteria for creating these categories: (a) each category should be clearly related to the phenomena studied, providing a specific perspective on how they are experienced; (b) the categories should have a hierarchical structure, progressing from simpler to more complex relationships; and (c) the category system should be parsimonious, explained with the fewest categories reasonably possible. Meeting these criteria makes the category system useful both theoretically and pedagogically [23]. These categories encompass two dimensions: the referential aspect and the structural aspect. The referential aspect pertains to the general meaning that individuals attribute to the phenomenon, reflecting the overarching interpretation that emerges in their experience. In contrast, the structural aspect focuses on how they organize and structure their understanding of the phenomenon. Both aspects are intrinsically linked and, together, facilitate a comprehensive representation of the phenomenon under investigation [24].
In this study, concepts are presented in descriptive categories according to Marton and Booth’s criteria. These categories are derived from the data obtained in the questionnaires and interviews, without attempting to force the data into pre-established categories. Each category highlights features that distinguish one concept from another and is organised hierarchically to reflect increasing levels of understanding. This hierarchy of descriptive categories shows the relationships between conceptions and provides a basis for teaching and assessment decisions [25].
3 Research and design framework
Our objective is to identify the different problem-solving approaches used by university students when applying the concept of circulation in Ampère-Maxwell’s law. A metacognitive question was given to 65 students from introductory electromagnetism courses at the University of the Basque Country (UPV/EHU) and the University of the Republic (UDELAR), representing all students enrolled in this subject at both institutions. The sample selection took into account that phenomenography aims to include individuals who have experienced the phenomenon in question; therefore, we chose students from the electromagnetism course to whom we had easy access. Additionally, to gain a deeper understanding of students’ thought processes, we conducted interviews with 12 students. We used convenience sampling to select students, ensuring data covered the full range of ways in which the phenomenon is experienced. We formed groups of 2 or 3 students, selecting them to represent a wide variety of academic levels and communication skills. This approach aimed to ensure the external validity of the results [24].
All students in the sample had completed the introductory electromagnetism course over 15 weeks, with 4 hours per week dedicated to theory and 2 hours to problem-solving. The content was based on the university physics textbooks by Walker, Resnick and Halliday [26], Tipler and Mosca [27], and Young and Freedman [28], with a week and a half devoted to the study of Ampère-Maxwell’s law. The courses were taught by experienced professors using a lecture-based methodology. In problem-solving sessions, the professor solved problems on the board in front of the group, using exercises from the end-of-chapter sections of the mentioned textbooks.
Once the question had been prepared, a preliminary test was carried out with students from the Electromagnetism course. It was confirmed that the students understood the way the proposed solution was written. In addition, the question was validated by consulting six lecturers with extensive experience in teaching electromagnetism. The validation of the interviews followed the same process of initial testing and peer validation.
@@ 3 Research and design framework
The question posed to the students shows a uniform electric field with magnitude increasing with time, confined to a cylindrical region of radius $R$, and states that the magnetic field is to be determined at a point $P$ located at a certain distance from the symmetry axis of the cylindrical region. The step-by-step solution proposed by an imaginary student, Agustina, to calculate the magnetic field as a function of the parameters of the problem is shown below. This question was selected because it presents a non-standard problem, unlike the examples found at the end of textbook chapters. It prompts students to think critically about how to solve the problem, evaluate the solution strategies provided, and, if necessary, propose alternative approaches. The proposed solution is correct, but it has the peculiarity that the closed curve chosen by the student exhibits a semicircle shape.
@@
    To calculate the magnetic field circulation, it is necessary to know the shape of the field lines in the context of the defined curve $C$, since its definition is a sum of elements $\vec{B}\cdot\overrightarrow{dl}$.
  3. 3.
    The Ampère-Maxwell law correlates the magnetic field circulation along the curve $C$ with the net current crossing a surface bounding it, without implying cause-effect relationships [29, 30].
The question was posed during the problem-solving class, after the students had already received lessons on the Ampère-Maxwell law. The answers were analysed by two of the researchers using phenomenographic methodology. At the beginning of the analysis, each researcher evaluated 20 written tests. A consensus was then reached on the categories of description that emerged from the students’ responses. Each researcher then analysed the remaining tests individually. At the end of this process, the results obtained were compared, and the reliability of the analysis was assessed using Cohen’s kappa coefficient, which yielded a value of 0.95, indicating significant agreement [31].
Each interview lasted approximately 20 minutes and in all the cases were conducted by the first author of this paper. The interviewer asked the students to discuss the question aloud and encouraged them to provide detailed explanations of their understanding using non-directive questions and follow-up questions such as “What do you mean by that?”, “Could you explain more?”, “Is there anything else you’d like to add about this step?”, “What do you think about the election of the curve $C$ to calculate the magnetic field?” (see A). All interviews were audio-recorded and transcribed. The extracts presented in the results section illustrate the sequence of interpretations and reasoning that characterise the processes of students’ development of situated cognition through problem solving. Fictitious names have been used in the presentation of results to ensure anonymity.
@@ 4 Results
According to the phenomenographic analysis, the emerging categories reflect varying degrees of understanding and reasoning of the students when reflecting metacognitively on the resolution of the problem (see Table 1). These categories are qualitatively differentiated according to the level of understanding shown when confronted with the question, standing out for both accuracy in the application of physical concepts and coherence in reasoning, and exhibit a hierarchy of understanding that reflects a progression in the ability to apply concepts coherently. Although the phenomenographic analysis does not focus on individual cases, the examples we provide as empirical justification for the categories necessarily pertain to responses from individual participants.
Table 1: Referential and structural aspects and the percentages obtained in each category.
Category Referential aspects Structural aspects %
A Agreement with the proposed solution. Represent the understanding of the magnetic field circulation and the magnetic field lines. Divide the curve into segments and analyse the direction of the magnetic field with respect to each segment to determine its contribution to the circulation. 15.4
B Disagreement with the calculation of the circulation. Refers to a lack of understanding of how certain segments of the curve contribute to the circulation of the field. They make incorrect assumptions about the contribution of each section of the curve to the flow of the field. 15.4
C Disagreement with the choice of the curve. Express the students’ idea that the selected curve should be symmetrical in order to be able to calculate the field. They use arguments that reflect an incomplete understanding of the role of symmetry in calculating the magnetic field from the AM law. 20.0
D Incoherent answers. Limited understanding of concepts, without articulation of a coherent strategy. 15.4
E No answer. Indicates lack of understanding or confidence to address the issue. 33.8
Category A (agreement with analysis), at the top of this hierarchy, groups together answers that correctly interpret the contribution of the magnetic field to the circulation along the selected curve, recognising that the magnetic field in the straight section of the curve does not contribute to the circulation. These students correctly apply the concept of magnetic field circulation, showing an understanding of the shape of magnetic field lines as well as the concept of circulation. The following is an example response in this category (fictitious names are used throughout the text).
@@
“Yes, because in $C_{1}$, $\vec{B}\cdot\overrightarrow{dl}=0$ since they are orthogonal throughout $C_{1}$, and in $C_{2}$, $\vec{B}\cdot$ $\overrightarrow{dl}=Bdl$ since they are collinear.” (Jimena, see Fig. 1)
Figure 2: Hand-drawn sketch given by Jimena.
@@
Jimena divides the closed curve into two sections, $C_{1}$ and $C_{2}$, and correctly recognises that the contribution to the magnetic field circulation along $C_{1}$ is null. This indicates that, although they do not directly represent the magnetic field along $C_{1}$, they understand that the magnetic field is perpendicular to each element of the curve. On the other hand, the student observes that the magnetic field is collinear with section $C_{2}$. In this way, they demonstrate an adequate understanding of the shape of the magnetic field lines and correctly interpret and apply the concept of magnetic field circulation.
Category B (disagreement with the calculation of the magnetic field circulation) is at an intermediate level. Students in this category show partial understanding as they consider that the straight part of the curve contributes to the magnetic field circulation. They are not able to see that the magnetic field is perpendicular to the straight line and that its contribution is zero. Here are two examples of answers in this category.
“dl is $(2+\pi)a$ and the result is $B=\frac{\mu_{0}\varepsilon_{0}}{(2+\pi)a}\frac{d\Phi_{\mathrm{E}}}{dt}$.”(Marcelo)
@@
“The induced field B is perpendicular to the field E at every point and its lines are closed. It is not perpendicular to the straight section of curve C, so it should have considered that section in the integral: $B(\pi a+2a)$.” (Juan)
@@
Marcelo believes that the magnetic field circulation should be calculated as the product of the magnetic field and the length of curve $C$. However, they do not recognise that the magnetic field along the straight section of curve $C$ does not contribute to the magnetic field circulation. On the other hand, Juan considers the direction of the magnetic field, acknowledging its perpendicularity to the electric field, but struggles to use this information effectively to calculate the magnetic field circulation, probably due to an incomplete understanding of circulation.
@@
Category C (disagreement with the choice of curve) reflects a lower level of understanding where students focus on the symmetry of the selected curve. In their answers, they express disagreement about the choice of the curve because, according to them, it should be symmetrical in order to be able to calculate the magnetic field. The following are two examples of responses in this category.
“No, because in my case the integration curve would be $2\pi a$ and not $\pi a$.” (María)
“I think it is wrong, because the whole surface is not affected by the electric field, you have to consider the whole circle.” (Ricardo)
From Maria’s answer, we can deduce that their disagreement with the solution of the question is due to the symmetry of the curve used, considering that a complete circle should be used. On the other hand, Ricardo argues that a complete circle should be used so that the entire variable electric field crosses the surface bounded by the curve.
@@
As we can see, Lucía clearly explains her reasoning. She states that the calculation of the magnetic field is only possible if the closed curve used is symmetrical. In another interview we identified similar reasoning to that of Lucía. In the following dialogue, three students exchange ideas about the choice of the curve C.
Florencia: To find the magnetic field at $P$, take a closed curve $C$ with the shape shown in the figure and apply the Ampère-Maxwell law. The law is correctly written.
Agustín: Yes. Since there’s no conduction current, ignore the current term.
Emilia: Now, regarding the choice of the curve, why take that curve? I mean, the worst curve in the world, right? What curve would we choose?
Florencia: We always go for things that are symmetric. Here, they’re half of symmetric.
Agustín: Sure, but what happens? You won’t be able to find the magnetic field using the curve you chose, and also because the magnetic field isn’t the same magnitude everywhere. The integral varies, and with that alone, you can’t calculate the field because you can calculate the circulation, but not the field. Do I make sense?
Interviewer: And you (Emilia and Florencia), what do you think about the choice of curve $C$ for calculating the magnetic field?
Emilia: It actually works, but you need to know how to do this integral (referring to the line integral of the magnetic field), which will be awful because the magnetic field varies and, besides, the curve isn’t symmetric.
Agustín: Hold on, what’s the issue? The problem we discussed several times—you’ll be able to calculate the circulation of the field; what you won’t be able to calculate is the field itself.
Emilia: Yes, but with our mathematical tools. Maybe she knows how to do that.
Agustín: What do we always do? What we do here is we say, the modulus of the field is constant here everywhere, so we take it outside and calculate a line integral. But what happens? Here you don’t calculate a line integral, you calculate a field integral of I don’t know what, I don’t know how much, and you calculate the circulation of the field, but you can’t…
Florencia: To solve the equation for $B$.
Agustín: Sure, you cannot get $B$.
Emilia: I agree.
When analysing the curve chosen to determine the magnetic field, Florencia mentions that they always select symmetrical curves, reflecting a strategy learned in class to solve problems of this type. Further on, when discussing the validity of the curve to find the magnetic field, the students argue that it is not possible to calculate it, as due to the shape of the curve, they cannot extract it from the integral. This suggests a limited understanding of the Ampère-Maxwell law, as they fail to interpret the magnetic field circulation as a sum of individual contributions and do not consider the specific shape of the magnetic field.
Finally, in another interview, we identified an exchange between two students that aligns with what Ricardo expressed. Below we present an extract from the interview in which the students exchange ideas about the shape of the closed curve.
@@
Category D (incoherent) includes answers without a logical structure or with ambiguous reasoning, where students show a limited and confused understanding of magnetic field circulation. At the lowest level of the hierarchy, Category E (no answer) represents no response, suggesting a lack of understanding or confidence in their ability to tackle the problem.
@@ 5 Discussion of results
Analysing the descriptive categories in relation to the research question, we find that they reveal different ways of understanding and applying magnetic field circulation, ranging from limited and incomplete interpretations to reflective thinking about the different processes needed to analyse and solve a problem.
@@
We found that around 15% adequately justify why they agree with all the steps in solving the question, showing an understanding of the application of magnetic field circulation in the context of Ampère-Maxwell’s law. We also observed that a similar percentage disagrees with the calculation of the magnetic field circulation, because they do not recognise that the magnetic field does not contribute to the circulation in the straight section. In this sense, we identified reasoning that could be related to the rote application of Ampère-Maxwell’s law, where students believe that the magnetic field circulation can always be expressed as the product between the magnetic field and the length of the closed curve. This type of reasoning has been documented in the literature in relation to Ampère’s law in different contexts [1, 2, 14].
@@
Furthermore, we found that in some cases students do not consider the shape of the magnetic field lines when applying Ampère-Maxwell’s law, while in others they do not use the magnetic field information adequately to calculate the magnetic field circulation, nor do they recognise that the magnetic field circulation is a sum (see for example the first interview in section 4). These findings converge with those already reported [14].
@@
We found that one in five students disagree with the choice of curve for calculating the circulation. Many of them believe that Ampère-Maxwell’s law can only be used to calculate magnetic fields if the curve used is symmetrical, which could be due to a functional fixation [32] associated with the most used procedure in class, which argues for symmetry to calculate magnetic fields using Ampère-Maxwell’s law.
@@
In addition, we found students who disagree with the shape of the curve used in the problem, considering that the variable electric field must cross the entire Amperian curve in order to be able to apply Ampère-Maxwell’s law and calculate the magnetic field. Our results suggest that a majority of students have an incomplete understanding of the Ampère-Maxwell law as they do not know that it correlates the magnetic field circulation with the net current that crosses a surface bounding it. This conceptual difficulty could be due to the belief that Ampère-Maxwell’s law establishes cause-effect relationships between the terms involved, in particular with the idea that currents (conduction and displacement) crossing a surface bounded by a curve must be the cause of the magnetic field computed in the line integral. This could lead them to argue that the curve must enclose the entire electric field. Another possible explanation is related to the belief that the magnetic field used in Ampère’s law is caused by enclosed currents [1, 2] and its origin could be in the way Maxwell’s equations are presented in most introductory physics textbooks, where they are described as cause-effect relationships between their different terms [16].